OpenStudy (anonymous):

What net force is required to bring a 1500 kg car to rest from a speed of 108 km/h in 15 s?

OpenStudy (anonymous):

108km/h/3600s=0.03km/s=30m/s 0=0.03km/s+15s*(a) -0.03/s=15s*(a) a=-0.03/15 a=0.002km/s^2=2m/s^2 also Where F is the force and m is the mass as well as a being the acceleration F=2m/s^2*1500kg=3000N

OpenStudy (anonymous):

Can someone verify this for me please

OpenStudy (anonymous):

I did the conversion from kilometers per hour to meters per second to comply with the rules of F=ma and then the rest consists of finding the acceleration and dividing with the correct mass to find the right answer .

OpenStudy (anonymous):

Overall my answer says I would need 3000N force to stop a truck weighing 1500kg from the velocity of 108km/h to rest within 15 seconds.

OpenStudy (anonymous):

Does that seem realistic overall or should I reconsider the solution thanks

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

U R RIGHT !

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Well I can't question your authority so yeah I think soXD

OpenStudy (anonymous):

Thanks guys:)

OpenStudy (anonymous):

you are right but it should be \[F=-3000N\] The force is acting in opposite direction to the vehicle, and this shows the retarding nature of the force the absolute value of the force is 3000N